@ Time stamp (to 1)6:34 The speaker converts x- solve 1.3 use the mean and standard deviation mentioned in the videol The speaker converts x#250 to z-1.67 using the z-score formula. Using the z-table used the NORMALCDFR.22,.2) what would your answer be taken 5 decimal places (the answer is slightly different)? back of book he determines 250) . 475 if you 7:04 The speaker determines P(2S0cx 35o) - .0475 by using the complement rule, etc. We understand that 250 and 350 are the boundaries. Again, using NORMALCDFR,2,2,2) what would your answer be taking 5 decimal places 2) 3) . 8:42 The speaker states he need to find-x based on given the Probabity. the bottom 5%based on 95% probablity. The speaker calculates 150.5. In my lecture video I refer to this as INVNORM(?,2,2) where we are doing the opposite of what we have already done (before we were given x and were inding probablity, Now we have the probablity e95%) and need to fnd x-in essence gong backwards, Take your INVNORM answer 2 decimal places. Email me: 1) P(X>250) answer should be 5 decimal places using NORMAL CDF 2) P(250<x<350)anser should be 5 decimal places using NORMALCO 3) Given the bottom 5% what is the value of x must take 2 decimal places?
Answer #1

(1) P(x > 250) = 0.04746

(2) P(250 < x < 350) = 0.90508

(3) z- score for bottom 5% is z = -1.6449

(There is no video available. But you can calculate this answer by doing the following calculation ...)

x = μ + z * σ (where μ = mean and σ = standard deviation)

x = μ - 1.6449 * σ